# Don’t Just Chase “Highlighted Tokens” in MLLMs: Revisiting Visual Holistic Context Retention

### v1

@@ 1 Introduction
As shown in Fig. 1, FastV [13] is an intuitive solution that ranks visual tokens based on attention distributions across different layers, and then prunes the bottom RR% of tokens based on the computational budget, thus reducing visual token redundancy. Subsequently, more work has followed this paradigm [88, 95, 4], designing different strategies to prune redundant visual tokens via cross-modal (i.e., text-vision) attention from LLMs. Besides, there are vision-centric pruning methods [75, 25, 91, 64] (e.g., FasterVLM [90]) that presume those visual tokens with low correlation to the [CLS] token in ViT [17], or those exhibit duplicated features tokens [20] to be redundant.
@@ 4.1.1 Fast Visual Context Refetching
Motivated by the attention sinks [93], and information loss during visual token pruning, we further propose visual context refetching to fast supplement the visual holistic context. Specifically, we treat pruned tokens as supplementary evidence, re-injecting them into the MLLM through Feed Forward Network (FFN) as “key-value memory” at the middle trigger layer. This refetch mechanism occurs when the model exhibits high uncertainty during inference, achieving effective and efficient visual information replenishment. Limited by space, the details can be found in Appendix E.
@@ 4.2 Theoretical Analysis
To further justify the trustworthiness of our proposed HoloV, we provide a theoretical analysis of it. Under Assumption 1, for any pruned token, there exists a retained token that is sufficiently close in the embedding space, with bounded context variance. By leveraging the Lipschitz continuity [8] of the transformer layer, we can bound the semantic difference between the outputs on the original and pruned token sets. The residual error introduced by the scoring threshold is also controlled. Combining these components, we obtain the stated upper bound. More details are in Appendix D.
@@ 5 Experiments
LLaVA-PruMerge (2024.5) 54.3 59.6 52.9 1632 71.3 67.9 70.6 54.3 50.1 91.4%
PDrop (2024.10) 57.1 63.2 56.8 1766 82.3 68.8 75.1 56.1 51.1 96.7%
@@
LLaVA-PruMerge (2024.5) 53.3 58.1 51.7 1554 67.2 67.1 68.8 54.3 50.3 89.4%
PDrop (2024.10) 56.0 61.1 56.6 1644 82.3 68.3 72.9 55.1 51.0 94.9%
@@
LLaVA-PruMerge (2024.5) 51.9 55.3 49.1 1549 65.3 68.1 67.4 54.0 50.1 87.7%
PDrop (2024.10) 41.9 33.3 50.5 1092 55.9 68.6 69.2 45.9 50.7 77.5%
@@ 5.1 Experimental Setup
Comparison methods. We compare our approach with several representative methods for accelerating multi-modal language models (MLLMs) via token reduction, including ToMe [11], FastV [13], SparseVLM [95], HiRED [4], LLaVA-PruMerge [64], PDrop [81], MustDrop [49], FasterVLM [90], and GlobalCom2[50]. These baselines employ diverse strategies such as token merging, attention-based pruning, adaptive allocation, and hierarchical retention to improve efficiency by reducing redundant tokens. Each method offers a unique perspective on balancing computational cost and model performance. More details of these baselines are provided in Appendix A.2.
@@ 5.3 HoloV with Higher Resolution
For further comprehensive evaluation, we also evaluated HoloV for LLaVA-NeXT on different benchmarks mentioned above, with comparison to current SOTA approaches. LLaVA-NeXT introduces a new image processing method, leading to dynamic lengths of visual embeddings for various image inputs. Thus, during the evaluation, 320 visual tokens has been kept (from up to 2880 raw tokens). As shown in Table 3, the evaluation results of all various benchmarks show that HoloV obtained the highest score on almost every track, and has an average of 95. 6%, much higher than the current SOTA of 93.3%. Besides, on video understanding benchmarks, HoloV maintains close to the original performance, significantly outperforming FasterVLM and FastV, as shown in Table 2. This demonstrates the value of HoloV when it comes to high-resolution visual input.
@@ 5.4 Efficiency Analysis
LLaVA-PruMerge (2024.5) 53.6 61.3 55.3 1534 60.8 66.4 69.7 50.6 54.0 85.6%
PDrop (2024.10) 56.4 63.4 56.2 1663 77.6 67.5 73.5 54.4 54.1 90.9%
@@ A.2 Backbones and Baselines
These methods collectively highlight diverse approaches to token reduction, ranging from attention-based pruning to adaptive merging, offering complementary solutions for accelerating MLLMs.
@@ A.3 Reproducibility
Appendix B Computational Complexity
To evaluate the computational complexity of MLLMs, it is essential to analyze their core components, including the self-attention mechanism and the feed-forward network (FFN). The total floating-point operations (FLOPs) required can be expressed as:
$\text{Total FLOPs}=T\times(4nd^{2}+2n^{2}d+2ndm),$ (7)
where $T$ denotes the number of transformer layers, $n$ is the sequence length, $d$ represents the hidden dimension size, and $m$ is the intermediate size of the FFN. This equation highlights the significant impact of sequence length $n$ on computational complexity. Notable, we follow FastV [13] to roughly estimate various token reduction baseline FLOPs. The FLOPs after token pruning can be represented as:
$\begin{split}\text{Post-Pruning FLOPs}\\ &\hskip-70.0001pt=L\times(4nd^{2}+2n^{2}d+2ndm)+{}\\ &\hskip-50.00008pt(T-L)\times(4\hat{n}d^{2}+2\hat{n}^{2}d+2\hat{n}dm),\end{split}$ (8)
where $L$ denotes the pruned layer, $\hat{n}$ represents token sequence length after pruning. The theoretical FLOPs reduction ratio related to visual tokens is computed as:
[…]
Appendix C More Sparsification Visualization
@@
C.1 MMBench Finegrained Results
@@
Appendix D Theoretical Analysis of HoloV
@@
Lemma D.1 (Token Coverage Guarantee)
@@
Theorem D.1 (Semantic Preservation)
@@
Proof D.2
Decompose the error into three components: 1) Geometric distortion: Bounded by Lemma D.1 2) Context variance: Controlled by $\sqrt{\delta}$ 3) Scoring residual:
@@
Appendix E Fast Visual Context Refetching
E.1 Preliminary: Reformulation of FFN
@@
E.2 FFN with Visual Context Refetching
@@
E.3 Further Efficiency Analysis
@@
Appendix F Impact Statement
@@ 8. Answer: [Yes]
  9. Justification: The discussion on the limitations of our work is stated in the paragraph F.
@@ 48. Answer: [Yes]
  49. Justification: The discussion on both potential positive societal impacts and negative societal impacts is stated in Appendix F.

### v2

@@ 1 Introduction
As shown in Fig. 1, FastV [13] is an intuitive solution that ranks visual tokens based on attention distributions across different layers, and then prunes the bottom RR% of tokens based on the computational budget, thus reducing visual token redundancy. Subsequently, more work has followed this paradigm [89, 96, 4], designing different strategies to prune redundant visual tokens via cross-modal (i.e., text-vision) attention from LLMs. Besides, there are vision-centric pruning methods [75, 25, 92, 64, 86] (e.g., FasterVLM [91]) that presume those visual tokens with low correlation to the [CLS] token in ViT [17], or those exhibit duplicated features tokens [20] to be redundant.
@@ 4.1.1 Fast Visual Context Refetching
Motivated by the attention sinks [94], and information loss during visual token pruning, we further propose visual context refetching to fast supplement the visual holistic context. Specifically, we treat pruned tokens as supplementary evidence, re-injecting them into the MLLM through Feed Forward Network (FFN) as “key-value memory” at the middle trigger layer. This refetch mechanism occurs when the model exhibits high uncertainty during inference, achieving effective and efficient visual information replenishment. Limited by space, the details can be found in Appendix D.
@@ 4.2 Theoretical Analysis
To further justify the trustworthiness of our proposed HoloV, we provide a theoretical analysis of it. Under Assumption 1, for any pruned token, there exists a retained token that is sufficiently close in the embedding space, with bounded context variance. By leveraging the Lipschitz continuity [8] of the transformer layer, we can bound the semantic difference between the outputs on the original and pruned token sets. The residual error introduced by the scoring threshold is also controlled. Combining these components, we obtain the stated upper bound. More details are in Appendix C.
@@ 5 Experiments
MustDrop (2024.11) 58.2 62.3 55.8 1787 82.6 69.2 76.0 56.5 51.4 97.2%
LLaVA-PruMerge (ICCV25) 54.3 59.6 52.9 1632 71.3 67.9 70.6 54.3 50.1 91.4%
PDrop (CVPR25) 57.1 63.2 56.8 1766 82.3 68.8 75.1 56.1 51.1 96.7%
@@
MustDrop (2024.11) 56.9 61.1 55.2 1745 78.7 68.5 74.6 56.3 52.1 95.7%
LLaVA-PruMerge (ICCV25) 53.3 58.1 51.7 1554 67.2 67.1 68.8 54.3 50.3 89.4%
PDrop (CVPR25) 56.0 61.1 56.6 1644 82.3 68.3 72.9 55.1 51.0 94.9%
@@
MustDrop (2024.11) 53.1 60.0 53.1 1612 68.0 63.4 69.3 54.2 51.2 90.1%
LLaVA-PruMerge (ICCV25) 51.9 55.3 49.1 1549 65.3 68.1 67.4 54.0 50.1 87.7%
PDrop (CVPR25) 41.9 33.3 50.5 1092 55.9 68.6 69.2 45.9 50.7 77.5%
@@ 5.1 Experimental Setup
Comparison methods. We compare our approach with several representative methods for accelerating multi-modal language models (MLLMs) via token reduction, including ToMe [11], FastV [13], SparseVLM [96], HiRED [4], LLaVA-PruMerge [64], PDrop [81], MustDrop [49], FasterVLM [91], GlobalCom2[50], VisionZip [86], DART [79]. These baselines employ diverse strategies such as token merging, attention-based pruning, adaptive allocation, and hierarchical retention to improve efficiency by reducing redundant tokens. Each method offers a unique perspective on balancing computational cost and model performance. More details of these baselines are provided in Appendix A.2.
@@ 5.3 HoloV with Higher Resolution
For further comprehensive evaluation, we also evaluated HoloV for LLaVA-NeXT on different benchmarks mentioned above, with comparison to current SOTA approaches. LLaVA-NeXT introduces a new image processing method, leading to dynamic lengths of visual embeddings for various image inputs. Thus, during the evaluation, 320 visual tokens has been kept (from up to 2880 raw tokens). As shown in Table 3, the evaluation results of all various benchmarks show that HoloV obtained the highest score on almost every track, and has an average of 95. 6%, much higher than the current SOTA of 93.3%.
@@
Besides, on video understanding benchmarks, HoloV maintains close to the original performance, significantly outperforming FasterVLM and FastV, as shown in Table 2. This demonstrates the value of HoloV when it comes to high-resolution visual input.
@@ 5.4 Efficiency Analysis
LLaVA-PruMerge (ICCV25) 53.6 61.3 55.3 1534 60.8 66.4 69.7 50.6 54.0 85.6%
PDrop (CVPR25) 56.4 63.4 56.2 1663 77.6 67.5 73.5 54.4 54.1 90.9%
@@ A.2 Backbones and Baselines
VisionZip [86] evaluates token importance via attention in the encoder and clustering retained tokens based on key similarity.
DART [79] introduces a duplication-aware token reduction method that selects a small subset of pivot tokens, calculates cosine similarity between pivot tokens and remaining tokens, retains those with the lowest duplication to pivots, achieving significant acceleration while maintaining performance and good compatibility with efficient attention operators. These methods collectively highlight diverse approaches to token reduction, ranging from attention-based pruning to adaptive merging, offering complementary solutions for accelerating MLLMs.
@@ A.3 Reproducibility
Appendix B More Sparsification Visualization
@@
B.1 MMBench Finegrained Results
@@
Appendix C Theoretical Analysis of HoloV
@@
Lemma C.1 (Token Coverage Guarantee)
@@
Theorem C.1 (Semantic Preservation)
@@
Proof C.2
Decompose the error into three components: 1) Geometric distortion: Bounded by Lemma C.1 2) Context variance: Controlled by $\sqrt{\delta}$ 3) Scoring residual:
@@
Appendix D Fast Visual Context Refetching
D.1 Preliminary: Reformulation of FFN
@@
D.2 FFN with Visual Context Refetching
@@
D.3 Further Efficiency Analysis
@@
Appendix E Impact Statement
@@ 8. Answer: [Yes]
  9. Justification: The discussion on the limitations of our work is stated in the paragraph E.
@@ 48. Answer: [Yes]
  49. Justification: The discussion on both potential positive societal impacts and negative societal impacts is stated in Appendix E.